taninverse root 3-cotinverse minus root 3
Answer by stanbon(75887)

taninverse root 3-cotinverse minus root 3
----
The angle whose tangent is sqrt(3) is pi pi/3 or 60 degrees
---
the angle whose cotangent is -sqrt(3)is (5/6)pi or 150 degrees
-----
Ans: pi/3 - (5/6)pi = (2/6 - 5/6)pi = -3/6 pi = -pi/2
or 60 = 150 = -90
============================
Cheers,
Stan H.
------------